a chemist has 6% and 15% solutions of acid his muvh of each solution should she mix to get 10 liters of a 9% solution
Answer by Alan3354(69443)

a chemist has 6% and 15% solutions of acid his muvh of each solution should she mix to get 10 liters of a 9% solution
----------
s = amount of 6%
f = amount of 15%
---
s + f = 10
6s + 15f = 10*9